# Flexible MOF Generation with Torsion-Aware Flow Matching

### v3

@@ 5.3 Preprocessing with canonicalization and MOF matching
To resolve this ambiguity, we introduce a canonicalization procedure that selects a unique rotation target for each symmetric building block. Given a clean conformation $\mathcal{C}_{1}=(A,X_{1})$ and a noised conformation $\mathcal{C}_{0}=(A,X_{0})$, we first identify the set of symmetry-preserving rotations of $\mathcal{C}_{0}$, defined as $\mathcal{R}:=\{\mathfrak{g}\in O(3)|\mathfrak{g}\cdot\mathcal{C}_{0}=\mathcal{C}_{0}\}\cap SO(3)$. We then apply each $\mathfrak{g}\in\mathcal{R}$ to $X_{0}$ to generate symmetrically equivalent conformations. Among these, we select the one with the lowest RMSD to $X_{1}$ and recompute the rotation between the two structures with Kabsch alignment [34] (Algorithm˜1). This yields the canonical rotation $\Delta q_{1\rightarrow 0}^{*}$, which we interpolate to obtain the final target rotation $\Delta q_{t\rightarrow 1}$ for any $t\in[0,1]$.
Canonicalization of torsion. MOF building blocks frequently contain $\pi$-symmetry bonds associated with distinct rotations that results in an equivalent conformation [17, 35]. Therefore, representing torsion as a relative rotation can lead to multiple valid targets for a single noisy conformation. To address this, we adopt a torsion representation that explicitly specifies the four atoms of the dihedral, with the neighboring atoms selected consistently using RDKit’s canonical atom rankings [14]. Concretely, given a rotatable bond between atoms $j$ and $k$, we select atoms $i$ and $l$ as the lowest-ranked neighbors of $j$ and $k$, respectively (Figure˜3(b)).
@@
MOF matching. We introduce a preprocessing step to reduce the distributional shift between the training and inference structures. Namely, during training, the model receives DFT-relaxed 3D conformations as input, while at inference, it receives coordinates initialized from the metal building block library and cheminformatics tools (see Section˜4.2). These small discrepancies in bond lengths and angles leads to degraded performance [16, 15]. To mitigate this, we propose MOF matching, a preprocessing procedure adapted from conformer matching from Jing et al. [16]. Specifically, MOF matching replaces each metal building block with its template structure from the library and each organic linker with an RDKit-generated structure whose torsion angles are optimized to closely match the original conformation. The resulting matched coordinates are used for training to ensure consistency with inference-time input. See Appendix˜E for details.
@@ 6 Experiments
Data preprocessing. We generally follow the preprocessing pipeline from prior work [11, 10]. Starting with the dataset from Boyd et al. [36], we apply metal-oxo decomposition algorithm from MOFid [37] and discard any structures containing more than 20 building blocks [11]. The resulting dataset is split into an 8:1:1 ratio for train/valid/test sets in the structure prediction task, and into a 9.5:0.5 train/valid split for MOF generation [11, 10]. Since the dataset [36] consists of hypothetical structures, we further filter out invalid MOFs using MOFChecker [38], and apply the MOF matching procedure to address distributional shift at test time, as discussed in Section˜5.3. Preprocessing details and dataset statistics are available in Appendix˜F.
@@ 6.1 MOF structure prediction
Baselines. We compare our model against both classic optimization- and learning-based approaches. For the classic baselines, we consider random search (RS) and the evolutionary algorithm (EA) implemented in CrySPY [39]. Among the learning-based baselines, we include DiffCSP [8], a general crystal structure prediction model, and MOFFlow [10], a MOF-specific structure prediction model. To ensure a fair comparison, we assume that no models have access to ground-truth local coordinates. Accordingly, we retrain MOFFlow with matched coordinates (Section˜5.3) and evaluate it using the initialized structures described in Section˜4.2.
Metrics. We use match rate (MR) and root mean square error (RMSE) computed with the StructureMatcher class from pymatgen [40]. StructureMatcher determines whether a predicted structure matches the reference based on specified tolerances. The RMSE is calculated only over the matched structures. We report results under two tolerance settings, $(0.5,0.3,10.0)$ and $(0.3,0.3,1.0)$, corresponding to site positions, fractional lengths, and lattice angle tolerances, respectively. The former is a standard threshold in CSP tasks [8, 9], while the latter is stricter for fine-grained evaluation. Note that while MOFFlow and MOFFlow-2 are trained on matched coordinates, we evaluate their predicted structures against ground-truth coordinates for fair comparison.
Results. Table˜2 shows that MOFFlow-2 outperforms all baselines across both tolerance settings. Consistent with prior work [10], we observe that the optimization-based methods (RS and EA) and the general CSP baseline (DiffCSP) achieve near-zero match rates, indicating their limited effectiveness on MOF structures. Compared to MOFFlow [10], our method achieves higher match rates and lower RMSE, underscoring the importance of explicitly modeling torsion angles for accurate structure prediction. We also provide visualizations of property distributions in Section˜G.1.
@@
RS EA DiffCSP [8] MOFFlow [10] MOFFlow-2
@@ 6.2 MOF generation
We evaluate the generative performance of MOFFlow-2 by (1) measuring validity, novelty, and uniqueness (VNU), and (2) comparing the distribution of MOF properties against that of the training set. We benchmark against MOFDiff [11], a coarse-grained diffusion model for MOF generation. Both models generate 10,000 samples using their respective pipelines. To isolate the effect of the generative model, no force-field relaxation is applied to the generated structures. Additional results on conditional generation and energy-level evaluation are provided in Section˜G.2 and Section˜G.3.
Metrics. A generated MOF is considered valid if it passes the MOFChecker test [38]. It is considered novel if its MOFid [37] does not appear in the training set. Uniqueness is computed as the proportion of distinct structures after removing duplicates from the generated set. We also report the percentage of novel building blocks (NBB), defined as the proportion of samples that are both valid and contain at least one building block not present in the training set. Additionally, we use Zeo++ [41] to evaluate MOF-specific properties, such as surface area and density.
@@ 7 Conclusion
We introduced MOFFlow-2, a two-stage generative model for MOF design and structure prediction. In the first stage, the building block generator designs MOF sequences in SMILES, which are initialized into 3D structures using a predefined metal library and RDKit [14]. In the second stage, the structure prediction module assembles the complete MOF by jointly predicting rotations, translations, torsions, and lattice parameters. Experimental results show that MOFFlow-2 outperforms existing models in both generative design and structure prediction.
@@ Acknowledgements
This work was supported by the NVIDIA Academic Grant Program. We thank NVIDIA for providing the GPU resources that enabled this research. It was also supported by the Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST)), National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2022-NR072184), GRDC(Global Research Development Center) Cooperative Hub Program through the National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2024-00436165), and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)).
We thank Vignesh Ram Somnath for his assistance in model scaling. We also thank Seonghyun Park, Hyomin Kim, and Junwoo Yoon for their contributions to improving the manuscript.
@@ Appendix A Implementation details for building block generator
We apply the SMILES tokenization regex from Schwaller et al. [28] to split the string into the following tokens.
@@
Batching. During training, we use dynamic batching [42] to prevent out-of-memory errors and reduce padding inefficiencies caused by highly variable sequence lengths. The key idea is to (1) limit each batch to a fixed maximum number of tokens and (2) group sequences of similar lengths. Specifically, we add the samples to a heap-based buffer sorted by sequence length. Once the buffer exceeds a predefined capacity, we add the longest sequences to the current batch until reaching the token limit, then start a new batch.
@@
Codebase. We implemented our sequence model with x-transformers [43]. We appreciate the author for the open-source implementation.
@@ C.1 Conditional flows and conditional vector fields
Torsions. Each torsion angle $\phi\in[-\pi,\pi)$ lies on a torus $\mathbb{T}$, where the exponential and logarithmic maps are defined as $\exp_{a}b=\operatorname{wrap}(a+b)$, $\log_{a}b=\operatorname{wrap}(b-a)$ with $\operatorname{wrap}(x)=(x+\pi)\operatorname{mod}{(2\pi)}-\pi$ [44]. The corresponding conditional flow and conditional vector field are:
@@
where $\exp:\mathfrak{so}(3)\rightarrow{}SO(3)$ and $\log:SO(3)\rightarrow{}\mathfrak{so}(3)$ are the exponential and logarithmic maps for $SO(3)$, respectively [45]. Since we target $\Delta q_{t\rightarrow{}1}:=q_{t}q_{1}^{\top}$ (Section˜5.1),
@@ C.2 Applying rotation, translation, and torsion
Torsions. We adapt the definition from Jing et al. [16] to apply torsion angles to MOF structures. Given coordinates $\bm{X}=[X^{(1)},\dots,X^{(M)}]$ and torsion angles $\bm{\phi}=[\phi^{(m)}\in\mathbb{T}^{P_{m}}]_{m=1}^{M}$, where $P_{m}$ is the number of rotatable bonds in the $m$-th building block, we update the coordinates around each rotatable bond $(j,k)$ with current angle $\phi$ and target angle $\phi^{\prime}$ as:
@@ D.1 Initialization module
    Ties between building blocks of the same type are resolved by sorting their centroid coordinates $(x,y,z)$ in ascending order of $x$, then $y$, then $z$ [46].
@@ D.2 Interaction module
The interaction module follows a Transformer encoder architecture [47] with root mean square layer normalization [RMSNorm; 48] (Figure˜2). Here, we provide further implementation details, including the model architecture (Algorithm˜4) and associated hyperparameters (Table˜5).
@@
1:# Construct edges and edge features [15]
@@ Appendix E MOF matching
We describe the MOF matching procedure, a preprocessing step used to reduce distributional shift between training and inference, as outlined in Section˜5.3. Given an MOF structure and its building blocks, we first determine whether each block contains metal elements to classify it as metal or organic. For metal blocks, we retrieve the corresponding template from the metal library and align it to the ground-truth structure. For organic blocks, we initialize the structure using RDKit, optimize torsion angles via differential evolution to minimize RMSD with the ground-truth [16], and then align the result. After processing all blocks, we compute the RMSD between the reconstructed and ground-truth structures. This procedure is repeated three times, and structures with final RMSD below 0.5Å are retained. The full algorithm for a single matching iteration is shown in Algorithm˜10.
@@
7:   # Organic building block [16]
@@ Appendix F Experimental details
Data preprocessing. We detail the data preprocessing steps described in Section˜6. First, we discard MOF structures with more than 20 building blocks since large structures that are too large may be difficult to synthesize [11]. Next, we extract key features from each structure, including Cartesian coordinates, RDKit-derived atomic features, Niggli-reduced cells, symmetrically equivalent coordinates for each building block (Section˜5.3), and canonical atoms defining the torsion angles (Section˜5.3). We then construct the metal library (Section˜4.2) and perform MOF matching (Appendix˜E) on the training dataset. Finally, since the dataset is synthetic [36], we filter out invalid MOFs using MOFChecker [38], which ensures the presence of key elements (e.g., C, H, and a metal), no atomic overlaps or coordination issues, sufficient porosity, and the absence of highly charged fragments or isolated molecules.
@@ F.1 Training details
Baselines. We follow Kim et al. [10] to RS and EA, using CrySPY[49] with CHGNet[50] for energy-based optimization. RS employs symmetry-based structure generation. For EA, we start with 5 random structures, select 4 parents via tournament selection, and generate offspring using 10 crossovers, 4 permutations, 2 strains, and 2 elites, iterating up to 20 generations.
For DiffCSP, we use a radius cutoff of 5Å and a batch size of 64. The model was trained for 500 epochs on an 80GB NVIDIA A100 GPU, taking approximately 5 days. Inference is performed with 1000 steps. All other settings follow the defaults in Jiao et al. [8]. FlowMM [9] is excluded from the baselines due to its high memory demands; based on our estimates, training it for 500 epochs would take over 30 days on an 80GB A100 GPU, which exceeds our resource constraints.
MOFFlow-2. We train our model on 8 80GB A100 GPUs for 200 epochs (about 4 days). To avoid out-of-memory issues caused by the large variation in MOF sizes, we use a dynamic batching strategy that limits each batch to a maximum of 1500 atoms. Specifically, the samples are first added to a buffer and, once the buffer exceeds a predefined capacity, they are added to the batch in a first-in-first-out manner until the maximum limit is reached. We use AdamW optimizer [51] with a learning rate of 1e-5, betas (0.9, 0.98), and no weight decay. Inference is performed in 50 steps.

### v4

@@ 5.3 Preprocessing with canonicalization and MOF matching
To resolve this ambiguity, we introduce a canonicalization procedure that selects a unique rotation target for each symmetric building block. Given a clean conformation $\mathcal{C}_{1}=(A,X_{1})$ and a noised conformation $\mathcal{C}_{0}=(A,X_{0})$, we first identify the set of symmetry-preserving rotations of $\mathcal{C}_{0}$, defined as $\mathcal{R}:=\{\mathfrak{g}\in O(3)|\mathfrak{g}\cdot\mathcal{C}_{0}=\mathcal{C}_{0}\}\cap SO(3)$. We then apply each $\mathfrak{g}\in\mathcal{R}$ to $X_{0}$ to generate symmetrically equivalent conformations. Among these, we select the one with the lowest RMSD to $X_{1}$ and recompute the rotation between the two structures with Kabsch alignment [kabsch1976solution] (Algorithm˜1). This yields the canonical rotation $\Delta q_{1\rightarrow 0}^{*}$, which we interpolate to obtain the final target rotation $\Delta q_{t\rightarrow 1}$ for any $t\in[0,1]$.
Canonicalization of torsion. MOF building blocks frequently contain $\pi$-symmetry bonds associated with distinct rotations that results in an equivalent conformation [jumper2021highly, zhang2023diffpack]. Therefore, representing torsion as a relative rotation can lead to multiple valid targets for a single noisy conformation. To address this, we adopt a torsion representation that explicitly specifies the four atoms of the dihedral, with the neighboring atoms selected consistently using RDKit’s canonical atom rankings [rdkit]. Concretely, given a rotatable bond between atoms $j$ and $k$, we select atoms $i$ and $l$ as the lowest-ranked neighbors of $j$ and $k$, respectively (Figure˜3(b)).
@@
MOF matching. We introduce a preprocessing step to reduce the distributional shift between the training and inference structures. Namely, during training, the model receives DFT-relaxed 3D conformations as input, while at inference, it receives coordinates initialized from the metal building block library and cheminformatics tools (see Section˜4.2). These small discrepancies in bond lengths and angles leads to degraded performance [jing2022torsional, corsodiffdock]. To mitigate this, we propose MOF matching, a preprocessing procedure adapted from conformer matching from jing2022torsional. Specifically, MOF matching replaces each metal building block with its template structure from the library and each organic linker with an RDKit-generated structure whose torsion angles are optimized to closely match the original conformation. The resulting matched coordinates are used for training to ensure consistency with inference-time input. See Appendix˜E for details.
@@ 6 Experiments
Data preprocessing. We generally follow the preprocessing pipeline from prior work [fu2023mofdiff, kim2025mofflow]. Starting with the dataset from boyd2019data, we apply metal-oxo decomposition algorithm from MOFid [bucior2019identification] and discard any structures containing more than 20 building blocks [fu2023mofdiff]. The resulting dataset is split into an 8:1:1 ratio for train/valid/test sets in the structure prediction task, and into a 9.5:0.5 train/valid split for MOF generation [fu2023mofdiff, kim2025mofflow]. Since the dataset [boyd2019data] consists of hypothetical structures, we further filter out invalid MOFs using MOFChecker [jablonka2023mofchecker], and apply the MOF matching procedure to address distributional shift at test time, as discussed in Section˜5.3. Preprocessing details and dataset statistics are available in Appendix˜F.
@@ 6.1 MOF structure prediction
Baselines. We compare our model against both classic optimization- and learning-based approaches. For the classic baselines, we consider random search (RS) and the evolutionary algorithm (EA) implemented in CrySPY [yamashita2018crystal]. Among the learning-based baselines, we include DiffCSP [jiao2024crystal], a general crystal structure prediction model, and MOFFlow [kim2025mofflow], a MOF-specific structure prediction model. To ensure a fair comparison, we assume that no models have access to ground-truth local coordinates. Accordingly, we retrain MOFFlow with matched coordinates (Section˜5.3) and evaluate it using the initialized structures described in Section˜4.2.
Metrics. We use match rate (MR) and root mean square error (RMSE) computed with the StructureMatcher class from pymatgen [ong2013pymatgen]. StructureMatcher determines whether a predicted structure matches the reference based on specified tolerances. The RMSE is calculated only over the matched structures. We report results under two tolerance settings, $(0.5,0.3,10.0)$ and $(0.3,0.3,1.0)$, corresponding to site positions, fractional lengths, and lattice angle tolerances, respectively. The former is a standard threshold in CSP tasks [jiao2024crystal, millerflowmm], while the latter is stricter for fine-grained evaluation. Note that while MOFFlow and MOFFlow-2 are trained on matched coordinates, we evaluate their predicted structures against ground-truth coordinates for fair comparison.
Results. Table˜2 shows that MOFFlow-2 outperforms all baselines across both tolerance settings. Consistent with prior work [kim2025mofflow], we observe that the optimization-based methods (RS and EA) and the general CSP baseline (DiffCSP) achieve near-zero match rates, indicating their limited effectiveness on MOF structures. Compared to MOFFlow [kim2025mofflow], our method achieves higher match rates and lower RMSE, underscoring the importance of explicitly modeling torsion angles for accurate structure prediction. We also provide visualizations of property distributions in Section˜G.1.
@@
RS EA DiffCSP [jiao2024crystal] MOFFlow [kim2025mofflow] MOFFlow-2
@@ 6.2 MOF generation
We evaluate the generative performance of MOFFlow-2 by (1) measuring validity, novelty, and uniqueness (VNU), and (2) comparing the distribution of MOF properties against that of the training set. We benchmark against MOFDiff [fu2023mofdiff], a coarse-grained diffusion model for MOF generation. Both models generate 10,000 samples using their respective pipelines. To isolate the effect of the generative model, no force-field relaxation is applied to the generated structures. Additional results on conditional generation and energy-level evaluation are provided in Section˜G.2 and Section˜G.3.
Metrics. A generated MOF is considered valid if it passes the MOFChecker test [jablonka2023mofchecker]. It is considered novel if its MOFid [bucior2019identification] does not appear in the training set. Uniqueness is computed as the proportion of distinct structures after removing duplicates from the generated set. We also report the percentage of novel building blocks (NBB), defined as the proportion of samples that are both valid and contain at least one building block not present in the training set. Additionally, we use Zeo++ [willems2012algorithms] to evaluate MOF-specific properties, such as surface area and density.
@@ 7 Conclusion
We introduced MOFFlow-2, a two-stage generative model for MOF design and structure prediction. In the first stage, the building block generator designs MOF sequences in SMILES, which are initialized into 3D structures using a predefined metal library and RDKit [rdkit]. In the second stage, the structure prediction module assembles the complete MOF by jointly predicting rotations, translations, torsions, and lattice parameters. Experimental results show that MOFFlow-2 outperforms existing models in both generative design and structure prediction.
@@ Acknowledgements
This work was supported by the NVIDIA Academic Grant Program. We thank NVIDIA for providing the GPU resources used in this work. This work was also supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education (RS-2025-25435147); Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST)); National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2022-NR072184), GRDC(Global Research Development Center) Cooperative Hub Program through the National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2024-00436165); and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)).
We thank Vignesh Ram Somnath for his help with model scaling. We also thank Seonghyun Park, Hyomin Kim, and Junwoo Yoon for their help improving the manuscript.
@@ Appendix A Implementation details for building block generator
We apply the SMILES tokenization regex from schwaller2019molecular to split the string into the following tokens.
@@
Batching. During training, we use dynamic batching [gardner2018allennlp] to prevent out-of-memory errors and reduce padding inefficiencies caused by highly variable sequence lengths. The key idea is to (1) limit each batch to a fixed maximum number of tokens and (2) group sequences of similar lengths. Specifically, we add the samples to a heap-based buffer sorted by sequence length. Once the buffer exceeds a predefined capacity, we add the longest sequences to the current batch until reaching the token limit, then start a new batch.
@@
Codebase. We implemented our sequence model with x-transformers [xtransformers2025]. We appreciate the author for the open-source implementation.
@@ C.1 Conditional flows and conditional vector fields
Torsions. Each torsion angle $\phi\in[-\pi,\pi)$ lies on a torus $\mathbb{T}$, where the exponential and logarithmic maps are defined as $\exp_{a}b=\operatorname{wrap}(a+b)$, $\log_{a}b=\operatorname{wrap}(b-a)$ with $\operatorname{wrap}(x)=(x+\pi)\operatorname{mod}{(2\pi)}-\pi$ [bose2024geometric]. The corresponding conditional flow and conditional vector field are:
@@
where $\exp:\mathfrak{so}(3)\rightarrow{}SO(3)$ and $\log:SO(3)\rightarrow{}\mathfrak{so}(3)$ are the exponential and logarithmic maps for $SO(3)$, respectively [yim2023se]. Since we target $\Delta q_{t\rightarrow{}1}:=q_{t}q_{1}^{\top}$ (Section˜5.1),
@@ C.2 Applying rotation, translation, and torsion
Torsions. We adapt the definition from jing2022torsional to apply torsion angles to MOF structures. Given coordinates $\bm{X}=[X^{(1)},\dots,X^{(M)}]$ and torsion angles $\bm{\phi}=[\phi^{(m)}\in\mathbb{T}^{P_{m}}]_{m=1}^{M}$, where $P_{m}$ is the number of rotatable bonds in the $m$-th building block, we update the coordinates around each rotatable bond $(j,k)$ with current angle $\phi$ and target angle $\phi^{\prime}$ as:
@@ D.1 Initialization module
    Ties between building blocks of the same type are resolved by sorting their centroid coordinates $(x,y,z)$ in ascending order of $x$, then $y$, then $z$ [bresson2025material].
@@ D.2 Interaction module
The interaction module follows a Transformer encoder architecture [vaswani2017attention] with root mean square layer normalization [RMSNorm; zhang2019root] (Figure˜2). Here, we provide further implementation details, including the model architecture (Algorithm˜4) and associated hyperparameters (Table˜5).
@@
1:# Construct edges and edge features [corsodiffdock]
@@ Appendix E MOF matching
We describe the MOF matching procedure, a preprocessing step used to reduce distributional shift between training and inference, as outlined in Section˜5.3. Given an MOF structure and its building blocks, we first determine whether each block contains metal elements to classify it as metal or organic. For metal blocks, we retrieve the corresponding template from the metal library and align it to the ground-truth structure. For organic blocks, we initialize the structure using RDKit, optimize torsion angles via differential evolution to minimize RMSD with the ground-truth [jing2022torsional], and then align the result. After processing all blocks, we compute the RMSD between the reconstructed and ground-truth structures. This procedure is repeated three times, and structures with final RMSD below 0.5Å are retained. The full algorithm for a single matching iteration is shown in Algorithm˜10.
@@
7:   # Organic building block [jing2022torsional]
@@ Appendix F Experimental details
Data preprocessing. We detail the data preprocessing steps described in Section˜6. First, we discard MOF structures with more than 20 building blocks since large structures that are too large may be difficult to synthesize [fu2023mofdiff]. Next, we extract key features from each structure, including Cartesian coordinates, RDKit-derived atomic features, Niggli-reduced cells, symmetrically equivalent coordinates for each building block (Section˜5.3), and canonical atoms defining the torsion angles (Section˜5.3). We then construct the metal library (Section˜4.2) and perform MOF matching (Appendix˜E) on the training dataset. Finally, since the dataset is synthetic [boyd2019data], we filter out invalid MOFs using MOFChecker [jablonka2023mofchecker], which ensures the presence of key elements (e.g., C, H, and a metal), no atomic overlaps or coordination issues, sufficient porosity, and the absence of highly charged fragments or isolated molecules.
@@ F.1 Training details
Baselines. We follow kim2025mofflow to RS and EA, using CrySPY[yamashita2021cryspy] with CHGNet[deng_2023_chgnet] for energy-based optimization. RS employs symmetry-based structure generation. For EA, we start with 5 random structures, select 4 parents via tournament selection, and generate offspring using 10 crossovers, 4 permutations, 2 strains, and 2 elites, iterating up to 20 generations.
For DiffCSP, we use a radius cutoff of 5Å and a batch size of 64. The model was trained for 500 epochs on an 80GB NVIDIA A100 GPU, taking approximately 5 days. Inference is performed with 1000 steps. All other settings follow the defaults in jiao2024crystal. FlowMM [millerflowmm] is excluded from the baselines due to its high memory demands; based on our estimates, training it for 500 epochs would take over 30 days on an 80GB A100 GPU, which exceeds our resource constraints.
MOFFlow-2. We train our model on 8 80GB A100 GPUs for 200 epochs (about 4 days). To avoid out-of-memory issues caused by the large variation in MOF sizes, we use a dynamic batching strategy that limits each batch to a maximum of 1500 atoms. Specifically, the samples are first added to a buffer and, once the buffer exceeds a predefined capacity, they are added to the batch in a first-in-first-out manner until the maximum limit is reached. We use AdamW optimizer [loshchilov2017decoupled] with a learning rate of 1e-5, betas (0.9, 0.98), and no weight decay. Inference is performed in 50 steps.